# Diffusion Model in Hyperspectral Image Processing and Analysis: A Review

### v1

@@ III Diffusion Model
Diffusion Model is an emerging generative model that has demonstrated powerful capabilities in the fields of image generation, denoising, and data enhancement in recent years. It gradually transforms noisy data into target data by modeling the diffusion process of data in time. The core idea of the diffusion model is to gradually transform the data into noise through the forward diffusion process, and then recover the original data from the noise through the reverse diffusion process. This approach performs well in high-dimensional data processing, image generation and denoising tasks, and especially has unique advantages in hyperspectral image processing.
III-A Theoretical Foundations of the Diffusion Model
@@
Diffusion model is divided into two processes: forward and backward diffusion diffusion; and finally training objectives are performed to minimize the difference between forward and backward diffusion processes. The basic diffusion models are usually divided into (1) denoising diffusion probabilistic model (DDPM), (2) score-based generative model (SGM), and stochastic differential equation-based model (SDE/NCSN). The basic principles of these three diffusion models are described below.
III-A1 Denoising Diffusion Probabilistic Model (DDPM)
@@
Denoising Diffusion Probabilistic Model (DDPM) [19, 22] utilize two Markov chains: a forward chain that perturbs the data into noise, and an inverse chain that converts the noise back into data. The former is usually manually designed to convert any data distribution into a simple prior distribution (e.g., a standard Gaussian distribution), while the latter Markov chain inverts the former by learning a transition kernel parameterized by a deep neural network.
@@
New data points are generated by first randomly sampling a vector from the previous distribution and ancestrally sampling it via an inverse Markov chain [23].
Figure 2: Diffusion processes in diffusion models
@@
Formally, given a data distribution $\mathbf{x}_{0}\sim q(\mathbf{x}_{0})$,, a forward Markov process generates a sequence of random variables $\mathbf{x}_{1},\mathbf{x}_{2}\ldots\mathbf{x}_{T}$ replaced by a transition kernel $q(x_{t}\mid{x_{t-1}})$. Using the probability chain rule and the Markov property, the joint distribution of $\mathbf{x}_{1},\mathbf{x}_{2}\ldots\mathbf{x}_{T}$ is later decomposed, conditional on $\mathbf{x}_{0}$, and denoted as $q(x_{1},\ldots{,x_{T}}\mid{x_{0}})$, which becomes:
@@
In DDPM, the transition kernel $q(x_{t}\mid{x_{t-1}})$ is drawn manually to incrementally transform the data distribution $q(\mathbf{x}_{0})$ into a processable prior distribution. A typical design of a transition kernel is a Gaussian perturbation, and the most common choice of transition kernel is:
@@
where $\beta_{t}\in(0,1)$ is the hyperparameter chosen prior to model training. This kernel is applied to simplify the discussion, although other types of kernels apply in the same vein. As observed by Sohl-Dickstein et al. 2015 [22] this Gaussian transition kernel allows us to marginalize the joint distribution in equation (1) to obtain an analytic form of $q(x_{t}\mid\mathbf{x}_{0})$ for all $t\in{0,1,\cdots,T}$. $\alpha_{t}:=1-\beta_{t}$ and $\bar{\alpha}_{t}{:}=\prod_{s=0}^{t}\alpha_{s}$, in particular:
@@
Intuitively, this forward process slowly injects noise into the data until all structure is lost. To generate new data samples, the DDPM first generates an unstructured noise vector from the prior distribution (which is usually easy to obtain), and then gradually removes the noise from it by running a learnable Markov chain in the opposite time direction. Specifically, the inverse Markov chain is parameterized by the prior distribution $p(\mathbf{x}_{T})=\mathcal{N}(\mathbf{x}_{T};0,\mathbf{I})$ and the learnable transition kernel $p_{\theta}(\mathbf{x}_{t-1}\mid\mathbf{x}_{t})$. We choose the prior distribution $p(\mathbf{x}_{T})=\mathcal{N}(\mathbf{x}_{T};0,\mathbf{I})$ because the forward process is constructed such that $q(\mathbf{x}_{T})\approx\mathcal{N}(\mathbf{x}_{T};0,\mathbf{I})$. The learnable transition kernel $p_{\theta}(\mathbf{x}_{t-1}\mid\mathbf{x}_{t})$ employs:
@@
where $\theta$ denotes the model parameters and the mean $\mu_{\theta}(\mathbf{x}_{t},t)$ and variance $\Sigma_{\theta}(\mathbf{x}_{t},t)$ are parameterized by the deep neural network.
The key to the success of the network training goal is to train the reverse process $p_{\theta}(\mathbf{x}_{0},\mathbf{x}_{1},\cdots,\mathbf{x}_{T})$ to match the actual time reversal of the forward process $q(\mathbf{x}_{0},\mathbf{x}_{1},\cdots,\mathbf{x}_{T})$. This is usually achieved by minimizing the KL scatter (Kullback-Leibler) between the two:
@@
where “const” is a constant that does not depend on the model parameters $\theta$. The goal of DDPM training is to maximize the variational lower bound or equivalently minimize the negative variational lower bound, and thus it can be efficiently estimated by Monte Carlo sampling [24] and efficiently optimized by stochastic optimization [25].
III-A2 Score-Based Generative Models (SGM/NCSN)
At the heart of Score-Based Generative Models (SGM/NCSN) [26, 20, 27] is the notion of a (Stein) score (aka fraction or score function) [28]. Given a probability density function p(x), the score function is defined as the gradient of the logarithmic probability density $\nabla_{\mathbf{x}}\log p(\mathbf{x})$. Unlike the Fisher score $\nabla_{\theta}\log p_{\theta}(\mathbf{x})$, which is commonly used in statistics, the Stein score considered here is a function of the data x and not of the model parameters $\theta$. It is a vector field pointing in the direction where the probability density function has the largest growth rate.
The key idea of Score-Based Generative Models (SGM) [20] is to perturb the data with a series of augmented Gaussian noises and jointly estimate the score function for all noisy data distributions by training a deep neural network model (called Noise Conditioned Scoring Network, NCSN, in [20]) conditional on the noise level. The samples are generated by linking the score functions with decreasing noise levels by using score-based sampling methods including Ronzmann Monte Carlo [29, 30, 31, 20, 21], stochastic differential equations [32, 21]], ordinary differential equations [33, 34, 35, 36, 21], and various combinations thereof [21]. In the formulation of the score-based generative model, training and sampling are completely decoupled, so that multiple sampling techniques can be used after estimating the score function.
@@ III-A2 Score-Based Generative Models (SGM/NCSN)
where $\lambda(t)$ is a positive weighting function, $\mathbf{x_{t}}$ is computed from $\varepsilon$ and $\mathbf{x}_{0}$, $\mathcal{U}[[1,T]]$ is a uniform distribution on the set ${1,2,\cdots,T}$, and $\varepsilon$ is a deep neural network with parameter $\theta$ for predicting the given $\mathbf{x}_{t}$ and t noise vectors. Comparing the equations, the training objectives of the DDPM and SGM are equivalent once $\epsilon_{\theta}(\mathbf{x},t)=-\sigma_{t}\mathbf{s}_{\theta}(\mathbf{x},t)$ is set.
III-A3 Diffusion model based on stochastic differential equations (Score-SDE)
When DDPM and SGM can be further generalized to the case of infinite time step or noise level, where the perturbation and denoising process is the solution of a stochastic differential equation (SDE), a formulation known as Score SDE [21], which is utilized for the noise perturbation and sample generation, and the denoising process requires estimation of a score function of the distribution of noisy data.
The SDE was used to score the data using a diffusion process controlled by the following stochastic differential equation (SDE) [21]:
@@ III-A3 Diffusion model based on stochastic differential equations (Score-SDE)
where f(x,t) and g(t) are the diffusion and drift functions of the SDE, and w is a standard Wiener process (aka Brownian motion) The forward processes in both the DDPM and the SGM are discretizations of this SDE. 2020 As demonstrated in the literature [21] by Song et al. the corresponding SDE for the DDPM is: $dx=-\frac{1}{2}\beta(t)xdt+\sqrt{\beta(t)}dw$. For the SGM the corresponding SDE is: $\mathrm{dx}=\sqrt{\frac{\mathrm{d}[\sigma(t)^{2}]}{\mathrm{d}t}}\mathrm{dw}$. where, $e\beta(\frac{t}{T})=T\beta_{t}$, $\sigma(\frac{t}{T})=\sigma_{t}$ when T tends to infinity.
Here $q_{t}({x})$ is used to denote the distribution of $\mathbf{x}_{t}$ in the forward process. For any diffusion process of the form of equation (14), Anderson proposed [42] to show that it can be inverted by solving the following reverse time SDE:
@@
where $\bar{\mathbf{W}}$ is the standard Wiener process when time flows backwards and dt denotes an infinitesimal negative time step. The solution trajectories of this reverse SDE have the same marginal density as those of the forward SDE, except that they evolve along opposite time directions [21]. Intuitively, the solution of a reverse-time SDE is a diffusion process that gradually converts noise into data. In addition, the literature presented by Song et al. in 2020 [21] proved the existence of an ordinary differential equation (ODE), the probabilistic flow ODE, whose trajectory has the same margin as the reverse time SDE. The probabilistic flow ODE is given by the following equation:
@@
Both the inverse time SDE and the probabilistic flow ODE allow sampling from the same data distribution because their trajectories have the same margins. If one knows the fractional function $\nabla_{\mathbf{x}}\log q_{t}(\mathbf{x})$ for each time step t, one can derive its inverse-time equations, which can then be used to generate samples by numerical techniques such as annealed Ronzwan dynamics [20], numerical SDE solvers [32, 21], numerical ODE solvers [33, 34, 21, 36, 43]], and predictor-corrector methods (a combination of MCMC and numerical ODE/SDE solver combination) [21]. As with the SGM, the transient fraction model $\mathbf{s}_{\theta}(\mathbf{x}_{t},t)$ is parameterized and the fraction function is estimated by generalizing the fraction matching objective in Equation (13) to continuous time, resulting in the following objective:
@@
where $\mathcal{U}[0,T]$ denotes the uniform distribution on $\mathbf{[}0,T]$.
The above is an introduction to the principles and formulas of the basic diffusion model, through their theoretical support, which makes the diffusion model work in the modern image field, solving certain limitations of the traditional model, but also brings defects. The study of Denoising Diffusion Probabilistic Models (DDPM) is the main focus in targeting the hyperspectral field, and it is also the most applied diffusion model in this field, the other two types of applications are relatively less, and they will also be introduced in the following sections.
III-B Evolution of the Diffusion model with different variants
Diffusion model started in 2015, Sohl-Dickstein et al. first proposed the theoretical framework of diffusion model, which combines nonequilibrium thermodynamics with generative model for the first time [22].In 2019, Song Yang et al. proposed Score-Based Generative Models (SGM/NCSN) [20], combining Score-Based Generative Models with Score Matching,which is a score-based generative model that generates by estimating the logarithmic gradient of the data distribution. Immediately following the optimization of the model in 2020, the team proposed NCSN++ (Noise-Conditioned Score Networks++) [27], which improves the network architecture (introducing residual connectivity and adaptive normalization), optimizes the noise scheduling strategy, and proposes a more stable training objective function. In the same year, Ho et al. proposed DDPM (Denoising Diffusion Probabilistic Models) [19], which used fixed noise scheduling and reparameterization techniques and became the basis of modern diffusion models. Later, in order to address the deficiencies of DDPM in terms of generation quality, training efficiency, and log-likelihood optimization, and so on, Improve-DDPM (IDDPM) [44] was proposed by Nichol et al. in 2021, which introduced learnable variance, cosine noise scheduling, and hybrid loss function, and accelerated the sampling strategy by time-step-importance sampling, which made the DDPM effectively improved , but the computation is complicated, the dataset has limitations, the generalization ability is insufficient, and the generation quality and speed belong to this and that.Song et al. proposed DDIM (Denoising Diffusion Implicit Models) [35] in the same year, which accelerates the sampling through non-Markov chains. Meanwhile, Rombach et al. proposed LDM (Latent Diffusion Model) [18], which shifts the diffusion process to latent space instead of operating directly in pixel space, dramatically reducing the computational cost and ensuring the generation quality. The enhancement for NCSN++ was also introduced in the same year with Score-SDE (based on stochastic differential equations) [21] as well as ODE (probabilistic streaming) [21], which optimize the sampling process with stochastic and constant differentiation and optimize the sampling process, respectively. Also in 2021, Dhariwal and Nichol proposed Classifier Guidance (Conditional Diffusion Model/CDM) [11], is to significantly improve the quality of generation by incorporating a classifier to guide the generation.In 2022, Rombach et al. published Stable Diffusion [18], which is based on the LDM model improved to enhance text-to-image generation. There are important variants [45, 46, 47, 48, 49, 50, 51, 52, 53, 54, 55, 56, 57] that have been improved since then, but do not constitute a “landmark” model. Below is a timeline of the evolution of the diffusion model and its main variants:

### v2

@@ 3 Diffusion Model
The Diffusion Model (DM) is an innovative generative model that has recently showcased remarkable capabilities in image generation, denoising, and data enhancement. Its core principle involves modeling a data diffusion process over time, progressively transforming noisy data into desired target data. Specifically, DMs operate by gradually converting original data into noise through a forward diffusion process, and then meticulously reconstructing the original data from this noise via a reverse diffusion process. This methodology proves highly effective for handling high-dimensional data, as well as for image generation and denoising tasks, demonstrating particular advantages in the realm of hyperspectral image processing.
3.1 Theoretical Foundations of the Diffusion Model
@@
The Diffusion Model operates through two distinct processes: forward diffusion and backward diffusion. The training objective is to minimize the discrepancy between these two processes. Fundamentally, diffusion models are typically categorized into three main types: (1) Denoising Diffusion Probabilistic Models (DDPM), (2) Score-Based Generative Models (SGM), and (3) Stochastic Differential Equation-based Models (SDE/NCSN). The foundational principles of each of these three diffusion models are described below.
3.1.1 Denoising Diffusion Probabilistic Model (DDPM)
@@
Denoising Diffusion Probabilistic Models (DDPMs) [19, 22] employ two Markov chains. A forward chain systematically introduces noise to transform the data into a simple prior distribution, typically a standard Gaussian. Conversely, an inverse chain learns to reverse this process, converting the noise back into data through a transition kernel parameterized by a deep neural network.
@@
New data points are generated by first randomly sampling a vector from the prior distribution, then using an inverse Markov chain to ancestrally sample it [23].
Figure 2: Diffusion processes in diffusion models [19]
@@
Formally, given a data distribution $\mathbf{x}_{0}\sim q(\mathbf{x}_{0})$, a forward Markov process generates a sequence of random variables, $\mathbf{x}_{1},\mathbf{x}_{2},\ldots,\mathbf{x}_{T}$, through a transition kernel $q(x_{t}\mid{x_{t-1}})$. By applying the probability chain rule and the Markov property, the joint distribution of these variables, conditional on $\mathbf{x}_{0}$, is then decomposed and denoted as $q(x_{1},\ldots,x_{T}\mid{x_{0}})$, which results in:
@@
In DDPMs, the transition kernel $q(x_{t}\mid{x_{t-1}})$ is specifically designed to gradually transform the data distribution $q(\mathbf{x}_{0})$ into a manageable prior distribution. A common and effective choice for this transition kernel is a Gaussian perturbation, typically defined as:
@@
where $\beta_{t}\in(0,1)$ is a hyperparameter set before model training. While other kernel types exist, this Gaussian transition kernel is used to simplify the discussion. As noted by Sohl-Dickstein et al. (2015) [22], this specific kernel allows for the marginalization of the joint distribution in equation (1), yielding an analytical form for $q(x_{t}\mid\mathbf{x}_{0})$ across all $t\in\{0,1,\cdots,T\}$. Specifically, with $\alpha_{t}:=1-\beta_{t}$ and $\bar{\alpha}_{t}{:}=\prod_{s=0}^{t}\alpha_{s}$:
@@
Intuitively, this forward process gradually introduces noise into the data until its original structure is completely lost. To generate new data samples, the DDPM first samples an unstructured noise vector from the prior distribution (which is typically straightforward to obtain). It then progressively removes this noise by running a learnable Markov chain in the reverse temporal direction. Specifically, this inverse Markov chain is parameterized by the prior distribution $p(\mathbf{x}_{T})=\mathcal{N}(\mathbf{x}_{T};0,\mathbf{I})$ and the learnable transition kernel $p_{\theta}(\mathbf{x}_{t-1}\mid\mathbf{x}_{t})$. We choose $p(\mathbf{x}_{T})=\mathcal{N}(\mathbf{x}_{T};0,\mathbf{I})$ for the prior because the forward process is designed such that $q(\mathbf{x}_{T})\approx\mathcal{N}(\mathbf{x}_{T};0,\mathbf{I})$. The learnable transition kernel $p_{\theta}(\mathbf{x}_{t-1}\mid\mathbf{x}_{t})$ employs:
@@
where $\theta$ represents the model’s parameters, while the mean $\mu_{\theta}(\mathbf{x}_{t},t)$ and variance $\Sigma_{\theta}(\mathbf{x}_{t},t)$ are parameterized by a deep neural network.
The key to successfully training the network lies in ensuring the reverse process $p_{\theta}(\mathbf{x}_{0},\mathbf{x}_{1},\cdots,\mathbf{x}_{T})$ accurately matches the true time reversal of the forward process $q(\mathbf{x}_{0},\mathbf{x}_{1},\cdots,\mathbf{x}_{T})$. This is typically achieved by minimizing the Kullback-Leibler (KL) divergence between the two:
@@
where "const" is a constant independent of the model parameters $\theta$. The objective of DDPM training is to maximize the variational lower bound, or equivalently, to minimize the negative variational lower bound. This can be efficiently estimated through Monte Carlo sampling [24] and optimized effectively via stochastic optimization [25].
3.1.2 Score-Based Generative Models (SGM/NCSN)
At the core of Score-Based Generative Models (SGM/NCSN) [26, 20, 27] lies the concept of a (Stein) score, also known as a fraction or score function [28]. For a given probability density function $p(\mathbf{x})$, the score function is defined as the gradient of the logarithmic probability density: $\nabla_{\mathbf{x}}\log p(\mathbf{x})$. Unlike the Fisher score, $\nabla_{\theta}\log p_{\theta}(\mathbf{x})$, which is commonly used in statistics and depends on the model parameters $\theta$, the Stein score is a function of the data $\mathbf{x}$ itself. Essentially, it’s a vector field that points in the direction of the steepest ascent of the probability density function.
The central idea behind Score-Based Generative Models (SGMs) [20] is to progressively add Gaussian noise to data and then collectively estimate the score function for all these noisy data distributions. This estimation is achieved by training a deep neural network, specifically known as a Noise Conditioned Scoring Network (NCSN) [20], which is conditioned on the noise level. Samples are then generated by linking these score functions across decreasing noise levels. This is done using various score-based sampling methods, including Ronzmann Monte Carlo [29, 30, 31, 20, 21], stochastic differential equations [32, 21], ordinary differential equations [33, 34, 35, 36, 21], and their combinations [21]. A key advantage of the SGM formulation is the complete decoupling of training and sampling, which allows for the use of multiple sampling techniques once the score function has been estimated.
@@ 3.1.2 Score-Based Generative Models (SGM/NCSN)
where $\lambda(t)$ represents a positive weighting function, while $\mathbf{x_{t}}$ is derived from $\varepsilon$ and $\mathbf{x}_{0}$. $\mathcal{U}[[1,T]]$ denotes a uniform distribution over the set $\{1,2,\ldots,T\}$, and $\varepsilon$ is a deep neural network with parameters $\theta$, tasked with predicting the noise vectors for a given $\mathbf{x}_{t}$ and $t$. Comparing the equations, the training objectives for DDPM and SGM become equivalent once $\epsilon_{\theta}(\mathbf{x},t)=-\sigma_{t}\mathbf{s}_{\theta}(\mathbf{x},t)$ is established.
3.1.3 Diffusion model based on stochastic differential equations (Score-SDE)
Both DDPM and SGM can be further generalized to an infinite time step or noise level. In this advanced formulation, known as Score SDE [21], the perturbation and denoising processes are governed by the solution of a stochastic differential equation (SDE). This approach is employed for both noise perturbation and sample generation, where the denoising process necessitates estimating a score function of the noisy data distribution.
The Stochastic Differential Equation (SDE) was employed to score the data using a diffusion process, which is governed by the following SDE [21]:
@@ 3.1.3 Diffusion model based on stochastic differential equations (Score-SDE)
where $f(\mathbf{x},t)$ and $g(t)$ represent the drift and diffusion functions of the SDE, respectively, while $w$ denotes a standard Wiener process (also known as Brownian motion). The forward processes in both the DDPM and SGM are essentially discretizations of this SDE. As demonstrated by Song et al. in 2020 [21], the corresponding SDE for the DDPM is given by $dx=-\frac{1}{2}\beta(t)xdt+\sqrt{\beta(t)}dw$. For the SGM, the corresponding SDE is $\mathrm{dx}=\sqrt{\frac{\mathrm{d}[\sigma(t)^{2}]}{\mathrm{d}t}}\mathrm{dw}$. This holds true as $T$ approaches infinity, where $e\beta(\frac{t}{T})=\beta_{t}T$ and $\sigma(\frac{t}{T})=\sigma_{t}$.
Here, $q_{t}({x})$ represents the distribution of $\mathbf{x}_{t}$ in the forward process. For any diffusion process structured like equation (14), Anderson [42] demonstrated that it can be inverted by solving the following reverse-time SDE:
@@
where $\bar{\mathbf{W}}$ represents the standard Wiener process as time flows backward, and $dt$ signifies an infinitesimal negative time step. The solution trajectories of this reverse SDE share the same marginal density as those of the forward SDE, but evolve in the opposite temporal direction [21]. Conceptually, the solution of a reverse-time SDE describes a diffusion process that progressively transforms noise into data. Furthermore, Song et al. (2020) [21] demonstrated the existence of an ordinary differential equation (ODE), known as the probability flow ODE, whose trajectory exhibits the same marginals as the reverse-time SDE. The probability flow ODE is defined by the following equation:
@@
Both the inverse time SDE and the probabilistic flow ODE allow for sampling from the same data distribution because their trajectories share identical marginals. If the fractional function $\nabla_{\mathbf{x}}\log q_{t}(\mathbf{x})$ is known for each time step $t$, its inverse-time equations can be derived. These equations can then be used to generate samples via numerical techniques such as annealed Ronzwan dynamics [20], numerical SDE solvers [32, 21], numerical ODE solvers [33, 34, 21, 36, 43], and predictor-corrector methods (a combination of MCMC and numerical ODE/SDE solvers) [21]. Similar to the SGM, the transient fractional model $\mathbf{s}_{\theta}(\mathbf{x}_{t},t)$ is parameterized, and the fractional function is estimated by generalizing the fractional matching objective in Equation (13) to continuous time, resulting in the following objective:
@@
where $\mathcal{U}[0,T]$ denotes the uniform distribution over the interval $[0,T]$.
The preceding discussion has introduced the fundamental principles and formulas of the basic diffusion model. This theoretical framework underpins the success of diffusion models in modern image processing, enabling them to overcome certain limitations of traditional models, albeit with their own inherent drawbacks. In the context of hyperspectral imaging, research primarily focuses on Denoising Diffusion Probabilistic Models (DDPMs), which are the most frequently applied diffusion models in this domain. While the other two types have seen comparatively less application, they will also be explored in subsequent sections.
3.2 Evolution of the Diffusion model with different variants
The genesis of Diffusion Models dates back to 2015, when Sohl-Dickstein et al. first introduced their theoretical framework, uniquely merging nonequilibrium thermodynamics with generative models [22]. In 2019, Song Yang et al. advanced the field with Score-Based Generative Models (SGM/NCSN) [20], which leverage score matching to generate data by estimating the logarithmic gradient of the data distribution. This was further refined in 2020 with NCSN++ (Noise-Conditioned Score Networks++) [27], optimizing the network architecture (through residual connections and adaptive normalization), noise scheduling, and introducing a more stable training objective. The same year saw Ho et al.’s seminal work on DDPM (Denoising Diffusion Probabilistic Models) [19], which, by utilizing fixed noise scheduling and reparameterization, established the foundation for modern diffusion models.
Addressing DDPM’s limitations in generation quality, training efficiency, and log-likelihood optimization, Nichol et al. introduced Improve-DDPM (IDDPM) [44] in 2021. This variant incorporated learnable variance, cosine noise scheduling, and a hybrid loss function, while also accelerating sampling through time-step-importance sampling. Despite these improvements, IDDPM still faced complexities in computation, dataset limitations, insufficient generalization, and trade-offs between generation quality and speed. Concurrently in 2021, Song et al. proposed DDIM (Denoising Diffusion Implicit Models) [35], accelerating sampling via non-Markovian chains. Significantly, Rombach et al. unveiled the Latent Diffusion Model (LDM) [18], which dramatically reduced computational cost and maintained generation quality by shifting the diffusion process to a latent space rather than operating directly on pixels. Further enhancements for NCSN++ emerged with Score-SDE (based on stochastic differential equations) [21] and ODE (probabilistic streaming) [21], both optimizing the sampling process through stochastic and constant differentiation, respectively. Also in 2021, Dhariwal and Nichol introduced Classifier Guidance (Conditional Diffusion Model/CDM) [11], significantly boosting generation quality by integrating a classifier to guide the process.In 2022, Rombach et al. released Stable Diffusion [18], an advancement based on the LDM model specifically enhanced for text-to-image generation.
Since then, numerous important variants have been developed [45, 46, 47, 48, 49, 50, 51, 52, 53, 54, 55, 56, 57], though none have yet reached the "landmark" status of the aforementioned models. A timeline illustrating the evolution of the diffusion model and its principal variants is provided below.